OpenStudy (anonymous):

1.) Solve by factoring: x^2+5x=6 x^2+5x=6 x^2+5x-6=0 (x+6)(x-1) CHECK x^2-1x+6x-6 x^2+5x-6 ANSWER: (x+6)(x-1) what did i do wrong?

OpenStudy (anonymous):

(x+6)(x-1) You did nothing wrong.

OpenStudy (anonymous):

my teacher said "Your first problem is not finished. You have to set each factor equal to 0 and solve for x."

OpenStudy (anonymous):

oooh nvm i got it lol

OpenStudy (anonymous):

Oh. In that case you need to discover when an x would result in an - -- or not.

OpenStudy (anonymous):

so i would just put (x+6)=0 x=6 (x-1)=0 x=-1

OpenStudy (anonymous):

The opposite. x+6=0 Therefore: x=-6 (x-1)=0 -> x=1

OpenStudy (anonymous):

ooooh k but -6+1= -5x not 5x

OpenStudy (anonymous):

@malical or @JakeV8

OpenStudy (anonymous):

I'm confused at your check. Based on your analysis, 6+6=0.

OpenStudy (anonymous):

ok....maybe we should just start the whole problem over lol

OpenStudy (anonymous):

i was using FOIL to check

OpenStudy (anonymous):

That just seems like a lot of typing. You need to review your answer again. I got x= -6,1 If you put -6 into the original equation you receive: 36-30=6 which checks out. Or at 1: 1+5=6 again.

OpenStudy (anonymous):

ooh i think your right